The area of a triangle is 165 square feet.If the height is 4 more than the base, find the length of the height and base of the triangle.
Answer by mananth(16946)

Area of triangle = 1/2*base * height.
Area = 165 sq.ft
let base be x
height = x+4
..
A= 1/2 * x*(x+4)
1/2 *(x^2+4x)=165
x^2+4x = 330
x^2+4x-330=0
solve for x using quadratic formula
b^2-4ac= 16+1320=346
(-4+sqrt(1336)) /2
=16.27 base
height = 20.27